The combined area of a square and a rectangle is 64 square centimeters. The width of the rectangle is 2 cm more than the length of a side of the square and the length of the rectangle is 2 cm more than its width. Find the dimensions of the square and the rectangle.
thanks a lot.. i'll really appreciate it if someone answered this one..
thanks again..

Answer by venugopalramana(3286)

The combined area of a square and a rectangle is 64 square centimeters. The width of the rectangle is 2 cm more than the length of a side of the square and the length of the rectangle is 2 cm more than its width. Find the dimensions of the square and the rectangle.
thanks a lot.. i'll really appreciate it if someone answered this one..
thanks again..
LET SIDE OF SQUARE =X CM.
2 CM MORE = X+2=WIDTH OF RECTANGLE
FURTHER 2 CM MORE =X+2+2=X+4=LENGTH OF RECTANGLE
AREA OF SQUARE = SIDE *SIDE =X*X=X^2
AREA OF RECTANGLE = LENGTH * WIDTH = (X+2)(X+4)=X^2+2X+4X+8=X^2+6X+8
TOTAL ARTEA OF SQUARE AND RECTANGLE = X^2+X^2+6X+8=2X^2+6X+8
=2(X^2+3X+4)=64...GIVEN
X^2+3X+4=64/2=32
X^2+3X+4-32=0
X^2+3X-28=0
X^2+7X-4X-28=0
X(X+7)-4(X+7)=0
(X+7)(X-4)=0
X-4=0...OR...X=4
SO SIDE OF SQUARE =4 CM.
WIDTH OF RECTANGLE =6 CM AND LENGTH OF RECTANGLE = 8 CM.